Meeting Notes for 14 Jan

Present: Bob, TomJ, Liz, Ken, Edit


  • Blessing:
    • Tom J described what we are still doing to finish step 1 of the blessfile recreation. We still have a few files left and need to decide what to do about them. Tom J is reviewing these.
    • For step 2 we need to decide whether we will re-bless the files. Tom J will forward the proposal again so we can discuss options next week.
  • Cpldfrequencies and shower analysis:
    • Tom will add a fix to split for the cpldfrequency calculation.
    • We will roll out that update to development and test the upload.
    • If we prove that this is a needed fix:
      • Need to rollout new version of split to production.
      • Need to update the cpldfrequency metadata for files between Oct 2011 and today.
      • Need to replace all threshold files for these splits.


  • Event Display:
    • Need to set up tools to get a snapshot of the server during a stress test to find out if opening many files causes an 'out of memory' error and the server gives up. Edit will work with this issue.


  • i2u2 hardware:
    • Systems at Argonne are working with the machines setup.
Main.EditPeronja - 2015-01-16
Topic revision: r1 - 2015-01-16, eperonja
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback